OpenStudy (anonymous):

Could someone explain how 4 + x^4 - 2 + (1/x^4) is equal to (x^2 + 1/(x^2))^2 please?

OpenStudy (anonymous):

is the first part really \[4+x^4-2\]??

OpenStudy (anonymous):

yes, + 1/(x^4)

OpenStudy (anonymous):

so first part is really \[2+x^4\] and you have \[2+x^4+\frac{1}{x^4}\]

OpenStudy (anonymous):

then to add you have a denominator of \[x^4\] so you need to write \[\frac{2x^4}{x^4}+\frac{x^8}{x^4}+\frac{1}{x^4}\] \[\frac{x^8+2x^4+1}{x^4}\] \[\frac{(x^4+1)^2}{x^4}\] or \[\left( \frac{x^4+1}{x^2}\right)^2\]

OpenStudy (anonymous):

Oh! Thanks! :)